OpenStudy (lilmike234):

let logba=3 logbc=2 logbd=5 what is the value of logb(A^5C^2/D^6) ? 1. 2 2. -15,378 3. -11 4. There isn't enough information to answer the question

OpenStudy (zzr0ck3r):

\(\log_b(\frac{A^5C^2}{D^6})=\log_b(A^5C^2)-\log_b(D^6)=\log_b(A^5)+\log_b(C^2)-\log_b(D^6)\\=5\log_b(A)+2\log_B(C)-6\log_b(D)\)

OpenStudy (lilmike234):

So my answer would be 3?

OpenStudy (zzr0ck3r):

how did you get that?

OpenStudy (lilmike234):

I plugged in 7 for the first log instead of 5. But I redid my calculations and got -11

OpenStudy (lilmike234):

Am I right ?

OpenStudy (zzr0ck3r):

correct